ALGEBRA1-Alternate Exam 12-continued Students Name Student Number On a 1287 mile long road trip in windy conditions, a familwith 17 and 18
Answer #1

17. Let the fuel economy of the vehicle in ideal (non-windy) conditions be f miles per gallon and let the influence of the wind be w miles per gallon.

Then f-w = 18…(1) and f+w = 26…(2).

On adding the 2 equations, we get 2f = 18+26 = 44 so that f = 44/2 = 22.

Now, on substituting f = 22 in the 2nd equation, we get 22+w = 26 so that w = 26-22 = 4.

Thus, the fuel economy of the vehicle in ideal (non-windy) conditions is expected to be 22 miles per gallon and the influence of the wind has an impact of 4 miles per gallon.

18. On the trip towards their desination, the vehicle consumed 1287/18 = 71.5 gallons of fuel and on the return trip, the vehicle consumed 1287/26 = 49.5 gallons of fuel. Therefore, during the return trip, the vehicle consumed 71.5+49.5 = 121 gallons of fuel.

In ideal (non-windy) conditions, the vehicle would consume 2*1287/22 = 117 gallons of fuel.

Thus, in windy conditions, the vehicle would consume 4 more gallons of fuel.